A tuning fork while vibrating with an air column at 51 degree celcius produces 4 beats and the same tuning fork produces 1 beats at 16 degree celcius. Find Frequency Of Tuning Fork. Ans 50Hz

Grade:

3 Answers

Ramesh V
14 years ago

when temp. decreases, speed of sound in air decreases, since no. of beats per second is less at lower tempertaure we coclude that frequency of aircolumn is higher than freq. of tuning fork.

let v be freq. of tuning fork, then at 51 degrees frq. of air column will be (n+4) where n is freq. of tuning fork

at 16 degrees , frewq. of air column will be (n+1)

v51/v16 = (n+4)*k / (n+1)*k
(n+4)/(n+1) = (273+51)1/2 / (273+16)1/2

on solving ofr n , we'll get n = 50 Hz

--

regards

Ramesh

Nehal Wani
14 years ago

Thank You very Much Sir.

Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

when temp. decreases, speed of sound in air decreases, since no. of beats per second is less at lower tempertaure we coclude that frequency of aircolumn is higher than freq. of tuning fork.

let v be freq. of tuning fork, then at 51 degrees frq. of air column will be (n+4) where n is freq. of tuning fork

at 16 degrees , frewq. of air column will be (n+1)v51/v16 = (n+4)*k / (n+1)*k
(n+4)/(n+1) = (273+51)1/2 / (273+16)1/2
on solving ofr n , we'll get n = 50 Hz

Thanks and Regards